if a multiple choice test consists of 5 questions each with 4 possible answers of which only one is correct,
a, in how many different ways can a student check off one answer to each question?
b, in how many ways can a student check off one answer to each question and get all the answers wrong?

Answer by stanbon(75887)

if a multiple choice test consists of 5 questions each with 4 possible answers of which only one is correct,
---------------------
a, in how many different ways can a student check off one answer to each question?
Ans: 4^5 = 2^10 = 1024
--------------------------------
b, in how many ways can a student check off one answer to each question and get all the answers wrong?
Ans: 3^5 = 243
====================
Cheers,
Stan H.
-------